An unbiased die is thrown find out the probability of getting

Prime number

Odd number

Even number


A multiple of 2&3

Number less than 7

A multiple of 4 ?


1
Expert's answer
2021-02-19T17:09:58-0500

Solution:


1) Prime numbers on a die are 1, 2, 3, 5; there are 4 in total:

"P(pr) = \frac{4}{6}=\frac{2}{3}=0.(6)"


2) Odd numbers on a die are 1, 3, 5; there are 3 in total:

"P(odd) = \frac{3}{6}=0.5"


3) Even numbers on a die are 2, 4, 6; there are 3 in total:

"P(even) = \frac{3}{6} = 0.5"


4) There is no 0 on a die, so:

"P(0) = 0"


5) Multiple of 2 and 3 is only 6:

"P(mul 2,3)= \frac{1}{6} = 0.1(6)"


6) All of the numbers on a die are less than 7:

"P(x<7) = 1"


7) The only multiple of 4 on a die is 4:

"P(4) = \frac{1}{6} = 0.1(6)"


Answer:

1) "0.(6)"

2) "0.5"

3) "0.5"

4)

5) "0.1(6)"

6) "1"

7) "0.1(6)"